# Quantum State Preparation with Optimal T-Count

### v1

@@ 1 Introduction
The Gottesman-Knill theorem [17] reveals a remarkable corner of Hilbert space occupied by Clifford circuits and stabilizer states — classically efficiently simulable but nevertheless equipped with genuinely quantum features such as many-body entanglement and superposition. Only non-Clifford operations can move us out of this corner, a requirement for universal quantum computation. A standard choice is to compile quantum circuits using the Clifford+${T}$ gate set, in which only the single-qubit ${T}=\mathsf{diag}(1,e^{i\pi/4})$ gate [^1] We use $\mathsf{diag}(c_{1},\ldots,c_{m})$ to denote the diagonal matrix with $c_{1},\ldots,c_{m}$ on the diagonal in order. is non-Clifford.
@@
In recent years there has been an interest in quantifying how much non-Cliffordness, aka magic, is required to implement quantum states and operations (see e.g., [46, 12, 4, 32]). This directly relates to the cost of fault-tolerant quantum computation based on 2D stabilizer error correcting codes such as the surface code, where the implementation of ${T}$ gates using magic state distillation is vastly more costly than the transversal implementation of Clifford gates [9]. [^2] A recent line of research [30, 21] shows that ${T}$ gates might not be as expensive as previously thought for certain architectures and fault tolerance schemes. A second reason to study quantum magic is its relevance to classical simulation of quantum computers: extensions of the Gottesman-Knill theorem give rise to classical simulation algorithms with a cost that scales polynomially in all parameters except the number of non-Clifford gates [12, 7, 4]. Another line of work has explored the role that magic plays in quantum many-body physics and phases of matter [33, 15].
@@
So how much magic is needed to make a quantum state, in the worst case? In particular, let us consider the number of ${T}$ gates, or ${T}$-count, required to $\varepsilon$-approximate an arbitrary $n$-qubit state to within $\varepsilon$-error in the $\ell_{2}$ norm. Here we consider unitary state preparation circuits composed of Clifford and T gates.
@@
For $n=1$, a remarkable series of works incorporated techniques from algebraic number theory to give a fairly complete answer to this question [27, 43, 41]. The algorithm of Ross and Selinger [41] computes a sequence of [^3] In this paper, asymptotic notations $O(\cdot),\Omega(\cdot)$ only hide absolute constants that do not depend on any parameter, and $\log(x)$ is the base 2 logarithm of $x$. $O(\log(1/\varepsilon))$ single-qubit ${H}$ and ${T}$ gates that $\varepsilon$-approximates a given single-qubit unitary $U$ (or prepares a $1$-qubit state $U|0\rangle$), matching information-theoretic lower bounds [24, 5]. These methods leverage beautiful structural properties of the group generated by single-qubit Clifford and ${T}$ gates established in [27] to outperform the well-known approximate synthesis technique based on the Solovay-Kitaev theorem [36].
@@
These methods can also be used to prepare an $n$-qubit state; one can first express it as $U|0^{n}\rangle$ where $U$ is a sequence of $O(2^{n})$ CNOT and single-qubit gates (see e.g., [3]), and then use the Ross-Selinger algorithm [41] to approximate each of the single-qubit gates as a sequence of ${H}$ and ${T}$ gates. In this way one can upper bound the number of ${T}$ gates for $n$-qubit state preparation as $O(2^{n}(n+\log(1/\varepsilon)))$. Note that this is also an upper bound on the total number of gates used. Surprisingly, [31, Section 3] showed that the ${T}$-count of $n$-qubit state synthesis can be improved by almost a square-root factor to
@@
The synthesis algorithm of [31] has two important ingredients. The first ingredient is a state preparation method proposed by Grover and Rudolph [18], which we now describe. For simplicity suppose the target state $|\psi\rangle=\sum_{x\in\{0,1\}^{n}}\alpha_{x}|x\rangle$ has non-negative real amplitudes $\alpha_{x}\geq 0$. Then, starting with $|0^{n}\rangle$, we first apply a single-qubit rotation on the first qubit to ensure it has the correct marginal, i.e., $|0\rangle\to\sqrt{\sum_{x:x_{1}=0}\alpha_{x}^{2}}|0\rangle+\sqrt{\sum_{x:x_{1}=1}\alpha_{x}^{2}}|1\rangle$. Then, controlled on the first qubit, we rotate the second qubit to have the correct marginal. We repeat this process until the $n$-th qubit. This reduces state preparation to the task of synthesizing $n$ multiply-controlled single-qubit unitaries, or (using Euler angles) $3n$ diagonal multiply-controlled single-qubit unitaries. The second ingredient is a method for approximating diagonal controlled single-qubit unitaries that starts with a $b=O(\log(n/\varepsilon))$-bit approximation of the rotation angle and then implements a controlled rotation corresponding to each bit. The technique used to implement the latter controlled rotations can be viewed as a generalization of the fact that an $n$-qubit Boolean diagonal unitary, with nonzero entries in ±$\pm 1$, can be implemented with ${T}$-count $\Theta\left(\sqrt{2^{n}}\right)$ [37, 42, 34, 31].
If we look beyond the headline square-root dependence $\sqrt{2^{n}}$, we can identify contributions to the ${T}$-count from each of these two ingredients that we might hope to improve: an overhead from the Grover-Rudolph method that scales with the number of qubits $n$ (under the square root in the first term of Equation 1) as well as an overhead that scales as $\log^{2}(n/\varepsilon)$ that arises from the implementing $b$ controlled single-qubit rotations, each to within $\varepsilon$-error.
In this work we show that indeed both of these overheads can be avoided. Our main result is the following improvement which gives an optimal upper bound on the ${T}$-count of an $n$-qubit state.
Theorem 1.1 (Quantum state preparation with optimal ${T}$-count).
Any $n$-qubit state can be prepared up to error $\varepsilon$ by a Clifford+${T}$ circuit starting with the all-zeros state using
@@
${T}$ gates and ancillas. Furthermore, no Clifford+${T}$ circuit (even with measurements and adaptivity) can use asymptotically fewer ${T}$ gates.
More precisely, for any $n$-qubit state $|\psi\rangle$, we show there is a Clifford+${T}$ circuit $U$ such that $U|0^{n}\rangle|0^{a}\rangle=|\widetilde{\psi}\rangle|0^{a}\rangle$, where $\||\psi\rangle-|\widetilde{\psi}\rangle\|\leq\varepsilon$. The ${T}$-count of $U$ and $a$, the number of ancillas used, are upper bounded by Equation 2.
The lower bound in Theorem 1.1, which we formally establish in Theorem 4.1, holds in an even stronger model that we call adaptive Clifford+${T}$ circuits (defined formally in Section 4). In this model, the algorithm may apply mid-circuit measurements and any future gates may depend on the outcomes of prior measurements. Such a circuit naturally outputs a mixed state, and even if we only require this state to be $\varepsilon$-close to the target state in trace distance, we show that the lower bound on ${T}$-count in Equation 2 still holds. The lower bound is established combining the $\Omega(\log(1/\varepsilon))$ lower bound from [5], which holds for adaptive Clifford+${T}$ circuits, and a slight generalization of the $\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ lower bound in [31], which only holds for unitary Clifford+${T}$ circuits. [^5] We thank Luke Schaeffer for confirming that their $\Omega\left(\sqrt{n2^{n}\log(1/\varepsilon)}\right)$ lower bound at the end of [31, Section 5] should actually be $\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$. Thus Theorem 1.1 represents the best of both worlds, with the upper bound being established in the weak model, but the lower bound being established in the strong model.
@@
For the first ingredient, we use a variant of recent methods from [25, 40] which reduce quantum state synthesis to the synthesis of $O(1)$ diagonal unitaries — far fewer than are needed in the Grover-Rudolph method. We start with a coarse approximation: we use just two Boolean diagonal unitaries to construct a state $|\phi\rangle$ that has a constant overlap with the target state $|\psi\rangle$ (i.e., $\varepsilon$ is a constant). Specifically, we show that there exist two diagonal unitaries $B_{1},B_{2}$ with diagonal entries in ±$\pm 1$, such that $|\phi\rangle:=B_{2}{H}^{\otimes n}B_{1}{H}^{\otimes n}|0^{n}\rangle$ has a constant overlap with $|\psi\rangle$. Starting from this coarse approximation, we use an idea from [40] to improve the accuracy to the desired value of $\varepsilon$. Since $|\phi\rangle$ already has a constant overlap with $|\psi\rangle$, intuitively their difference $|\psi\rangle-|\phi\rangle$ should have a small length. Then our goal becomes to construct $|\psi\rangle-|\phi\rangle$, for which we again have a coarse approximation. Iterating this, we obtain $|\psi\rangle$ in the limit, where the error decays exponentially: $|\psi\rangle$ is $\varepsilon$-close to
@@
For the second ingredient, we give a new method to synthesize diagonal unitaries with optimal ${T}$-count.
Theorem 1.2 (Diagonal unitary synthesis with optimal ${T}$-count).
Any diagonal unitary on $n$ qubits can be implemented up to error $\varepsilon$ by a Clifford+${T}$ circuit using
@@
${T}$ gates and ancillas. Furthermore, no Clifford+${T}$ circuit (even with measurements and adaptivity) can use asymptotically fewer ${T}$ gates.
Here, the distance between two unitaries is the operator norm (aka spectral norm) of their difference, i.e., $U$ implements $V$ to error $\varepsilon$ if $\|U-V\|=\varepsilon$. The lower bound in Theorem 1.2, which is formally proved in Theorem 4.2, holds in the stronger model of adaptive Clifford+${T}$ circuits, as before. The lower bound is proved by a reduction to the previous lower bound; see Section 4 for details.
To prove the upper bound from Theorem 1.2, we first note that an $n$-qubit diagonal unitary $D$ can be regarded as a single-qubit diagonal unitary $G_{y}$ controlled on $n-1$ qubits $|y\rangle$. Now we use the fact that any single-qubit unitary $G_{y}$ can be approximated by a sequence of $O(\log(1/\varepsilon))$ Hadamard and ${T}$ gates (using no ancillas) [27, 43, 28, 41]. To synthesize $D$, we start by applying a Boolean oracle $B\colon|y\rangle|z\rangle|0\rangle\to|y\rangle|z\rangle|s_{y}\rangle$, where $y\in\{0,1\}^{n-1},z\in\{0,1\}$, and $s_{y}$ is a binary string of length $O(\log(1/\varepsilon))$ that describes a sequence of ${H}$ and ${T}$ gates that $\varepsilon$-approximates $G_{y}$. This Boolean oracle $B$ can be implemented with ${T}$-count $O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ [34, 31]. After this step, we apply a sequence of controlled-Hadamard and controlled-${T}$ gates on the $n$-th qubit, each controlled on a bit of $|s_{y}\rangle$. The total ${T}$-cost of this last step is $O(\log(1/\varepsilon))$, proportional to the length of $s_{y}$. Finally, we uncompute the Boolean oracle $B$, returning the third register to the all-zeros state.
We note that it is possible to directly improve on the state synthesis technique from [31] using Theorem 1.2; this gives a better ${T}$-count than Equation 1, but still worse than the optimal bound Equation 2. See Appendix B for details.
@@ 1.1 Applications
There are some immediate gains by using our bounds in place of those from [31] in known applications. For instance, we obtain slight improvements to the state-of-the-art lower bound on the stabilizer rank of magic states, due to Mehraban and Tahmasbi [35]. [^6] In particular, we can shave off two factors of $\log m$ to get a slightly improved bound for the approximate stabilizer rank $\chi_{\delta}(|{T}\rangle^{\otimes m})=\Omega\left(\frac{m^{2}}{\log^{2}m}\right)$ for $\delta=\Omega(1)$. Our results also improve some related lower bounds concerning decompositions of Boolean functions as linear combination of quadratic functions [16]; see [35] for details. We can also obtain improved bounds for “partial unitary synthesis” where the goal is to implement the first $K$ columns of an $n$-qubit unitary. Here state preparation corresponds to the case of $K=1$ and unitary synthesis is the case of $K=2^{n}$. Using Householder reflections [23], this can be reduced to the task of preparing up to $K$ different $n$-qubit states [26, 31]. Via this reduction and with our Theorem 1.1, we obtain a ${T}$-count of $O\left(K\sqrt{2^{n}\log(K/\varepsilon)}+K\log(K/\varepsilon)\right)$, which similarly shaves polylogrithmic factors from the bounds in [31].
@@
Results from [43, 5] show that synthesizing a single-qubit unitary has ${T}$-count $\Theta(\log(1/\varepsilon))$. What is the cost of implementing $m$ single-qubit unitaries $U_{1}\otimes\cdots\otimes U_{m}$ to error $\varepsilon$? [^7] Here $\varepsilon$ is the operator-norm error in approximating the entire unitary $U_{1}\otimes\cdots\otimes U_{m}$. The most natural bound would be $O(m\log(m/\varepsilon))$ by implementing each up to $\varepsilon/m$ error. Surprisingly, a corollary of our previous result implies that the total cost remains the same $\log(1/\varepsilon)$ even if $m$ has slightly superconstant dependence on $1/\varepsilon$.
@@
Let $m=O(\log\log(1/\varepsilon))$ be an integer. A tensor product of $m$ (potentially different) single-qubit unitaries $U_{1}\otimes\cdots\otimes U_{m}$ can be implemented up to error $\varepsilon$ by a Clifford+${T}$ circuit using $O\left(\log(1/\varepsilon)\right)$ ${T}$ gates and ancillary qubits.
Theorem 1.3 is a simple consequence of Theorem 1.2. To see this, note that we can use the Euler angle decomposition to express each single-qubit unitary $U_{j}$ as a sequence of Hadamards and at most $3$ diagonal unitaries, i.e, $U_{j}=D_{3,j}{H}D_{2,j}{H}D_{1,j}$, where $D_{i,j}$ are diagonal single-qubit unitaries for $i\in\{1,2,3\}$ and $j\in[m]$. This reduces our synthesis task to the case in which all the single-qubit unitaries are diagonal; clearly it is enough to synthesize $D_{i,1}\otimes D_{i,2}\otimes\cdots\otimes D_{i,m}$ for $i=1,2,3$. But now we can apply Theorem 1.2 which implies that any $\log\log(1/\varepsilon)$-qubit diagonal unitary can be implemented with ${T}$-count $O(\log(1/\varepsilon))$.
@@
The batched synthesis result (Theorem 1.3) focused on implementing $m$ different single-qubit unitaries. It is natural to ask what happens if we want to prepare $m$ identical single-qubit unitaries. This is called “mass production” [45, 44, 29] in circuit synthesis and lives in a richer family of direct sum problems; see the survey [38] for details. As an application of Theorem 1.2, we prove the following optimal ${T}$-count for mass production of a single-qubit unitary.
@@
Let $U$ be an arbitrary single-qubit unitary. Then $U^{\otimes m}$ can be implemented up to error $\varepsilon$ by a Clifford+${T}$ circuit using $O\left(m+\log(1/\varepsilon)\right)$ ${T}$ gates and ancillary qubits. Furthermore, no Clifford+${T}$ circuit (even with measurements and adaptivity) can use asymptotically fewer ${T}$ gates.
The bound in Theorem 1.4 is indeed better than the one in Theorem 1.3 if the goal is to implement the same unitary several times: Theorem 1.4 shows that using $O(\log(1/\varepsilon))$ ${T}$ gates, we can prepare $m=O(\log(1/\varepsilon))$ copies of a given single-qubit unitary $U$, in contrast to Theorem 1.3, which says we can implement $m=O(\log\log(1/\varepsilon))$ (potentially different) unitaries.
Theorem 1.4 is also a corollary of Theorem 1.2. Using the Euler angle decomposition, we can reduce to the case where $U$ is diagonal, say, $U=\mathsf{diag}(1,e^{i\theta})$. Then $U^{\otimes m}|x\rangle=e^{i\theta\cdot h_{x}}|x\rangle$, where $h_{x}\in\left\{0,1,\ldots,m\right\}$ is the Hamming weight of $x\in\{0,1\}^{m}$. Therefore, we can first compute the Hamming weight into a second register containing $L=\lceil\log m\rceil$ qubits $|x\rangle\to|x\rangle|h_{x}\rangle$, then apply the corresponding phase change $|h_{x}\rangle\to e^{i\theta\cdot h_{x}}|h_{x}\rangle$. The first step can be done with $O(m)$ complexity in a standard way (see Fact 2.6 for details). For the second step, since $h_{x}$ has $L$ bits, the phase change is a diagonal unitary on $L=\left\lceil\log m\right\rceil$ qubits, which can be implemented with ${T}$-count $O\left(\sqrt{m\log(1/\varepsilon)}+\log(1/\varepsilon)\right)\leq O\left(m+\log(1/\varepsilon)\right)$ by Theorem 1.2.
Note that one can implement a similar but sub-optimal strategy without the use of Theorem 1.2, see, e.g., [5, Section 3.2]. Indeed, one can implement the second step as $R(\theta)\otimes R(2\theta)\otimes\ldots R(2^{L}\theta)|h_{x}\rangle=e^{i\theta\cdot h_{x}}|h_{x}\rangle$ where $R(\alpha)\equiv\mathsf{diag}(1,e^{i\alpha})$ and then use the Ross-Selinger algorithm to approximate each single-qubit gate $R(2^{k}\theta)$ to within error $\varepsilon/L$. This strategy gives an overall ${T}$-count of

### v2

@@ 1 Introduction
The Gottesman-Knill theorem [17] reveals a remarkable corner of Hilbert space occupied by Clifford circuits and stabilizer states — classically efficiently simulable but nevertheless equipped with genuinely quantum features such as many-body entanglement and superposition. Only non-Clifford operations can move us out of this corner, a requirement for universal quantum computation. A standard choice is to compile quantum circuits using the Clifford+T{T} gate set, in which only the single-qubit ${T}=\mathsf{diag}(1,e^{i\pi/4})$ gate [^1] We use $\mathsf{diag}(c_{1},\ldots,c_{m})$ to denote the diagonal matrix with $c_{1},\ldots,c_{m}$ on the diagonal in order. is non-Clifford.
@@
In recent years there has been an interest in quantifying how much non-Cliffordness, aka magic, is required to implement quantum states and operations (see e.g., [47, 12, 4, 32]). This directly relates to the cost of fault-tolerant quantum computation based on 2D stabilizer error correcting codes such as the surface code, where the implementation of T{T} gates using magic state distillation is vastly more costly than the transversal implementation of Clifford gates [9]. [^2] A recent line of research [30, 21] shows that T{T} gates might not be as expensive as previously thought for certain architectures and fault tolerance schemes. A second reason to study quantum magic is its relevance to classical simulation of quantum computers: extensions of the Gottesman-Knill theorem give rise to classical simulation algorithms with a cost that scales polynomially in all parameters except the number of non-Clifford gates [12, 7, 4]. Another line of work has explored the role that magic plays in quantum many-body physics and phases of matter [33, 15].
@@
So how much magic is needed to make a quantum state, in the worst case? In particular, let us consider the number of T{T} gates, or T{T}-count, required to $\varepsilon$-approximate an arbitrary $n$-qubit state to within $\varepsilon$-error in the $\ell_{2}$ norm. Here we consider unitary state preparation circuits composed of Clifford and T gates.
@@
For $n=1$, a remarkable series of works incorporated techniques from algebraic number theory to give a fairly complete answer to this question [27, 43, 41]. The algorithm of Ross and Selinger [41] computes a sequence of [^3] In this paper, asymptotic notations $O(\cdot),\Omega(\cdot)$ only hide absolute constants that do not depend on any parameter, and $\log(x)$ is the base 2 logarithm of $x$. $O(\log(1/\varepsilon))$ single-qubit H{H} and T{T} gates that $\varepsilon$-approximates a given single-qubit unitary $U$ (or prepares a $1$-qubit state $U|0\rangle$), matching information-theoretic lower bounds [24, 5]. These methods leverage beautiful structural properties of the group generated by single-qubit Clifford and T{T} gates established in [27] to outperform the well-known approximate synthesis technique based on the Solovay-Kitaev theorem [36].
@@
These methods can also be used to prepare an $n$-qubit state; one can first express it as $U|0^{n}\rangle$ where $U$ is a sequence of $O(2^{n})$ CNOT and single-qubit gates (see e.g., [3]), and then use the Ross-Selinger algorithm [41] to approximate each of the single-qubit gates as a sequence of H{H} and T{T} gates. In this way one can upper bound the number of T{T} gates for $n$-qubit state preparation as $O(2^{n}(n+\log(1/\varepsilon)))$. Note that this is also an upper bound on the total number of gates used. Surprisingly, [31, Section 3] showed that the T{T}-count of $n$-qubit state synthesis can be improved by almost a square-root factor to
@@
The synthesis algorithm of [31] has two important ingredients. The first ingredient is a state preparation method proposed by Grover and Rudolph [18], which we now describe. For simplicity suppose the target state $|\psi\rangle=\sum_{x\in\{0,1\}^{n}}\alpha_{x}|x\rangle$ has non-negative real amplitudes $\alpha_{x}\geq 0$. Then, starting with $|0^{n}\rangle$, we first apply a single-qubit rotation on the first qubit to ensure it has the correct marginal, i.e., $|0\rangle\to\sqrt{\sum_{x:x_{1}=0}\alpha_{x}^{2}}|0\rangle+\sqrt{\sum_{x:x_{1}=1}\alpha_{x}^{2}}|1\rangle$. Then, controlled on the first qubit, we rotate the second qubit to have the correct marginal. We repeat this process until the $n$-th qubit. This reduces state preparation to the task of synthesizing $n$ multiply-controlled single-qubit unitaries, or (using Euler angles) $3n$ diagonal multiply-controlled single-qubit unitaries. The second ingredient is a method for approximating diagonal controlled single-qubit unitaries that starts with a $b=O(\log(n/\varepsilon))$-bit approximation of the rotation angle and then implements a controlled rotation corresponding to each bit. The technique used to implement the latter controlled rotations can be viewed as a generalization of the fact that an $n$-qubit Boolean diagonal unitary, with nonzero entries in $\pm 1$, can be implemented with T{T}-count $\Theta\left(\sqrt{2^{n}}\right)$ [37, 42, 34, 31].
If we look beyond the headline square-root dependence $\sqrt{2^{n}}$, we can identify contributions to the T{T}-count from each of these two ingredients that we might hope to improve: an overhead from the Grover-Rudolph method that scales with the number of qubits $n$ (under the square root in the first term of Equation 1) as well as an overhead that scales as $\log^{2}(n/\varepsilon)$ that arises from implementing $b$ controlled single-qubit rotations, each to within $\varepsilon$-error.
In this work we show that indeed both of these overheads can be avoided. Our main result is the following improvement which gives an optimal upper bound on the T{T}-count of an $n$-qubit state.
Theorem 1.1 (Quantum state preparation with optimal T{T}-count).
Any $n$-qubit state can be prepared up to error $\varepsilon$ by a Clifford+T{T} circuit starting with the all-zeros state using
@@
T{T} gates and ancillas. [^5] This construction uses $O(2^{n}\log(1/\varepsilon))$ Clifford gates. Furthermore, no Clifford+T{T} circuit (even with measurements and adaptivity) can use asymptotically fewer T{T} gates.
More precisely, for any $n$-qubit state $|\psi\rangle$, we show there is a Clifford+T{T} circuit $U$ such that $U|0^{n}\rangle|0^{a}\rangle=|\widetilde{\psi}\rangle|0^{a}\rangle$, where $\||\psi\rangle-|\widetilde{\psi}\rangle\|\leq\varepsilon$. The T{T}-count of $U$ and $a$, the number of ancillas used, are upper bounded by Equation 2.
The lower bound in Theorem 1.1, which we formally establish in Theorem 4.1, holds in an even stronger model that we call adaptive Clifford+T{T} circuits (defined formally in Section 4). In this model, the algorithm may apply mid-circuit measurements and any future gates may depend on the outcomes of prior measurements. Such a circuit naturally outputs a mixed state, and even if we only require this state to be $\varepsilon$-close to the target state in trace distance, we show that the lower bound on T{T}-count in Equation 2 still holds. The lower bound is established combining the $\Omega(\log(1/\varepsilon))$ lower bound from [5], which holds for adaptive Clifford+T{T} circuits, and a slight generalization of the $\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ lower bound in [31], which only holds for unitary Clifford+T{T} circuits. [^6] We thank Luke Schaeffer for confirming that their $\Omega\left(\sqrt{n2^{n}\log(1/\varepsilon)}\right)$ lower bound at the end of [31, Section 5] should actually be $\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$. Thus Theorem 1.1 represents the best of both worlds, with the upper bound being established in the weak model, but the lower bound being established in the strong model.
@@
For the first ingredient, we use a variant of recent methods from [25, 40] which reduce quantum state synthesis to the synthesis of $O(1)$ diagonal unitaries — far fewer than are needed in the Grover-Rudolph method. We start with a coarse approximation: we use just two Boolean diagonal unitaries to construct a state $|\phi\rangle$ that has a constant overlap with the target state $|\psi\rangle$ (i.e., $\varepsilon$ is a constant). Specifically, we show that there exist two diagonal unitaries $B_{1},B_{2}$ with diagonal entries in $\pm 1$, such that $|\phi\rangle:=B_{2}{H}^{\otimes n}B_{1}{H}^{\otimes n}|0^{n}\rangle$ has a constant overlap with $|\psi\rangle$. Starting from this coarse approximation, we use an idea from [40] to improve the accuracy to the desired value of $\varepsilon$. Since $|\phi\rangle$ already has a constant overlap with $|\psi\rangle$, intuitively their difference $|\psi\rangle-|\phi\rangle$ should have a small length. Then our goal becomes to construct $|\psi\rangle-|\phi\rangle$, for which we again have a coarse approximation. Iterating this, we obtain $|\psi\rangle$ in the limit, where the error decays exponentially: $|\psi\rangle$ is $\varepsilon$-close to
@@
For the second ingredient, we give a new method to synthesize diagonal unitaries with optimal T{T}-count.
Theorem 1.2 (Diagonal unitary synthesis with optimal T{T}-count).
Any diagonal unitary on $n$ qubits can be implemented up to error $\varepsilon$ by a Clifford+T{T} circuit using
@@
T{T} gates and ancillas. [^7] This construction uses $O(2^{n}\log(1/\varepsilon))$ Clifford gates. Furthermore, no Clifford+T{T} circuit (even with measurements and adaptivity) can use asymptotically fewer T{T} gates.
Here, the distance between two unitaries is the operator norm (aka spectral norm) of their difference, i.e., $U$ implements $V$ to error $\varepsilon$ if $\|U-V\|=\varepsilon$. The lower bound in Theorem 1.2, which is formally proved in Theorem 4.2, holds in the stronger model of adaptive Clifford+T{T} circuits, as before. The lower bound is proved by a reduction to the previous lower bound; see Section 4 for details.
To prove the upper bound from Theorem 1.2, we first note that an $n$-qubit diagonal unitary $D$ can be regarded as a single-qubit diagonal unitary $G_{y}$ controlled on n1n-1 qubits $|y\rangle$. Now we use the fact that any single-qubit unitary $G_{y}$ can be approximated by a sequence of $O(\log(1/\varepsilon))$ Hadamard and T{T} gates (using no ancillas) [27, 43, 28, 41]. To synthesize $D$, we start by applying a Boolean oracle $B\colon|y\rangle|z\rangle|0\rangle\to|y\rangle|z\rangle|s_{y}\rangle$, where $y\in\{0,1\}^{n-1},z\in\{0,1\}$, and $s_{y}$ is a binary string of length $O(\log(1/\varepsilon))$ that describes a sequence of H{H} and T{T} gates that $\varepsilon$-approximates $G_{y}$. This Boolean oracle $B$ can be implemented with T{T}-count $O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ [34, 31]. After this step, we apply a sequence of controlled-Hadamard and controlled-T{T} gates on the $n$-th qubit, each controlled on a bit of $|s_{y}\rangle$. The total T{T}-cost of this last step is $O(\log(1/\varepsilon))$, proportional to the length of $s_{y}$. Finally, we uncompute the Boolean oracle $B$, returning the third register to the all-zeros state.
We note that it is possible to directly improve on the state synthesis technique from [31] using Theorem 1.2; this gives a better T{T}-count than Equation 1, but still worse than the optimal bound Equation 2. See Appendix B for details.
@@ 1.1 Applications
There are some immediate gains by using our bounds in place of those from [31] in known applications. For instance, we obtain slight improvements to the state-of-the-art lower bound on the stabilizer rank of magic states, due to Mehraban and Tahmasbi [35]. [^8] In particular, we can shave off two factors of $\log m$ to get a slightly improved bound for the approximate stabilizer rank $\chi_{\delta}(|{T}\rangle^{\otimes m})=\Omega\left(\frac{m^{2}}{\log^{2}m}\right)$ for $\delta=\Omega(1)$. Our results also improve some related lower bounds concerning decompositions of Boolean functions as linear combination of quadratic functions [16]; see [35] for details. We can also obtain improved bounds for “partial unitary synthesis” where the goal is to implement the first $K$ columns of an $n$-qubit unitary. Here state preparation corresponds to the case of $K=1$ and unitary synthesis is the case of $K=2^{n}$. Using Householder reflections [23], this can be reduced to the task of preparing up to $K$ different $n$-qubit states [26, 31]. Via this reduction and with our Theorem 1.1, we obtain a T{T}-count of $O\left(K\sqrt{2^{n}\log(K/\varepsilon)}+K\log(K/\varepsilon)\right)$, which similarly shaves polylogarithmic factors from the bounds in [31].
@@
Results from [43, 5] show that synthesizing a single-qubit unitary has T{T}-count $\Theta(\log(1/\varepsilon))$. What is the cost of implementing $m$ single-qubit unitaries $U_{1}\otimes\cdots\otimes U_{m}$ to error $\varepsilon$? [^9] Here $\varepsilon$ is the operator-norm error in approximating the entire unitary $U_{1}\otimes\cdots\otimes U_{m}$. The most natural bound would be $O(m\log(m/\varepsilon))$ by implementing each up to $\varepsilon/m$ error. Surprisingly, a corollary of our previous result implies that the total cost remains the same $\log(1/\varepsilon)$ even if $m$ has slightly superconstant dependence on $1/\varepsilon$.
@@
Let $m=O(\log\log(1/\varepsilon))$ be an integer. A tensor product of $m$ (potentially different) single-qubit unitaries $U_{1}\otimes\cdots\otimes U_{m}$ can be implemented up to error $\varepsilon$ by a Clifford+T{T} circuit using $O\left(\log(1/\varepsilon)\right)$ T{T} gates and ancillary qubits. [^10] This construction uses $O(\log(1/\varepsilon))$ Clifford gates.
Theorem 1.3 is a simple consequence of Theorem 1.2. To see this, note that we can use the Euler angle decomposition to express each single-qubit unitary $U_{j}$ as a sequence of Hadamards and at most $3$ diagonal unitaries, i.e, $U_{j}=D_{3,j}{H}D_{2,j}{H}D_{1,j}$, where $D_{i,j}$ are diagonal single-qubit unitaries for $i\in\{1,2,3\}$ and $j\in[m]$. This reduces our synthesis task to the case in which all the single-qubit unitaries are diagonal; clearly it is enough to synthesize $D_{i,1}\otimes D_{i,2}\otimes\cdots\otimes D_{i,m}$ for $i=1,2,3$. But now we can apply Theorem 1.2 which implies that any $\log\log(1/\varepsilon)$-qubit diagonal unitary can be implemented with T{T}-count $O(\log(1/\varepsilon))$.
@@
The batched synthesis result (Theorem 1.3) focused on implementing $m$ different single-qubit unitaries. It is natural to ask what happens if we want to prepare $m$ identical single-qubit unitaries. This is called “mass production” [46, 45, 29] in circuit synthesis and lives in a richer family of direct sum problems; see the survey [38] for details. As an application of Theorem 1.2, we prove the following optimal T{T}-count for mass production of a single-qubit unitary.
@@
Let $U$ be an arbitrary single-qubit unitary. Then $U^{\otimes m}$ can be implemented up to error $\varepsilon$ by a Clifford+T{T} circuit using $O\left(m+\log(1/\varepsilon)\right)$ T{T} gates and ancillary qubits. [^11] This construction uses $O(m\log(1/\varepsilon))$ Clifford gates. Furthermore, no Clifford+T{T} circuit (even with measurements and adaptivity) can use asymptotically fewer T{T} gates.
The bound in Theorem 1.4 is indeed better than the one in Theorem 1.3 if the goal is to implement the same unitary several times: Theorem 1.4 shows that using $O(\log(1/\varepsilon))$ T{T} gates, we can prepare $m=O(\log(1/\varepsilon))$ copies of a given single-qubit unitary $U$, in contrast to Theorem 1.3, which says we can implement $m=O(\log\log(1/\varepsilon))$ (potentially different) unitaries.
Theorem 1.4 is also a corollary of Theorem 1.2. Using the Euler angle decomposition, we can reduce to the case where $U$ is diagonal, say, $U=\mathsf{diag}(1,e^{i\theta})$. Then $U^{\otimes m}|x\rangle=e^{i\theta\cdot h_{x}}|x\rangle$, where $h_{x}\in\left\{0,1,\ldots,m\right\}$ is the Hamming weight of $x\in\{0,1\}^{m}$. Therefore, we can first compute the Hamming weight into a second register containing $L=\lceil\log m\rceil$ qubits $|x\rangle\to|x\rangle|h_{x}\rangle$, then apply the corresponding phase change $|h_{x}\rangle\to e^{i\theta\cdot h_{x}}|h_{x}\rangle$. The first step can be done with $O(m)$ complexity in a standard way (see Fact 2.6 for details). For the second step, since $h_{x}$ has $L$ bits, the phase change is a diagonal unitary on $L=\left\lceil\log m\right\rceil$ qubits, which can be implemented with T{T}-count $O\left(\sqrt{m\log(1/\varepsilon)}+\log(1/\varepsilon)\right)\leq O\left(m+\log(1/\varepsilon)\right)$ by Theorem 1.2.
Note that one can implement a similar but sub-optimal strategy without the use of Theorem 1.2, see, e.g., [5, Section 3.2]. Indeed, one can implement the second step as $R(\theta)\otimes R(2\theta)\otimes\ldots R(2^{L}\theta)|h_{x}\rangle=e^{i\theta\cdot h_{x}}|h_{x}\rangle$ where $R(\alpha)\equiv\mathsf{diag}(1,e^{i\alpha})$ and then use the Ross-Selinger algorithm to approximate each single-qubit gate $R(2^{k}\theta)$ to within error $\varepsilon/L$. This strategy gives an overall T{T}-count of